OpenStudy (anonymous):

can you solve ln(3x-5)=0

OpenStudy (anonymous):

Yes. You should know that ln (1) = 0, So you want to know when ln(3x-5) = 0, Well, that will occur when 3x - 5 = 1 or x= 2

OpenStudy (anonymous):

Alternate Solution: Ln is base e So write ln(3x - 5) = 0 as e^0 = 3x-5 or 1 = 3x-5, or x = 2 e^0 = 1

OpenStudy (anonymous):

ok, thank you! :)

OpenStudy (anonymous):

Welcome.